THE NEAR INFRARED TRANSITION OF COPPER CHLORIDE BY INTRACAVITY LASER SPEC-TROSCOPY

\underline{J.J. O’BRIEN}, Department of Chemistry, University of Missouri, St. Louis, MO 63121-4499; L.C. O’BRIEN, Department of Chemistry, Southern Illinois University, Edwardsville, IL 62026-1652.

The near infrared electronic transition of CuCl occurring in the region of 13,500 cm^{−1}has been recorded by intracavity laser absorption spectroscopy. The CuCl molecules were produced in a copper hollow cathode operating with an applied potential of 700 V, using 1.4 torr argon and a trace amount of carbon tetrachloride. The near infrared transition previously was recorded by Balfour and Ram,^{a} and it was assigned as emission from a lowlying ^{3}Σ^{+} state to the X^{1}Σ^{+} state. Recently, this transition has received considerable attention because of high level ab initio calculations ^{b} that question whether this transition involves the X^{1}Σ^{+} ground state.

aW.J. Balfour and R.S. Ram, J. Phys. B 17, L19 (1984).

bC. Sousa, W.A. de Jong, R. Broer and W.C. Nieuwpoort, Mol. Phys. 92, 677 (1997).